OpenStudy (anonymous):

cos2 theta=cos^2(theta-1/2)

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

I need to solve it but im not even sure where to start

OpenStudy (anonymous):

okay... cos(2theta) or cos^2(theta)

OpenStudy (anonymous):

my book just says "solve the equation" so i guess both

OpenStudy (anonymous):

first term is cos^2(theta) or cos(2theta)

OpenStudy (anonymous):

cos(2theta)

OpenStudy (anonymous):

okay.... cos^2+sin^2=1

OpenStudy (anonymous):

cos^2(theta)=1-sin^2(theta)

OpenStudy (anonymous):

and then what do i do with the right side of the equation?

OpenStudy (anonymous):

sin^2(theta)=1/2-1/2cos(2theta)

OpenStudy (anonymous):

cos^2(theta)=1-(1/2-1/cos2(theta) cos^2(theta)=1-1/2+1/2cos(2theta) cos^2(theta)=(1+cos(2theta)/2